OpenStudy (anonymous):

The sum of the first 30 terms of the sequence an = 6n + 5 is _____. *** PLease explain cuz iheartfood is confused :(

OpenStudy (anonymous):

Hi!!! THanks for coming to help again :)

OpenStudy (anonymous):

the sum formular is \[S _{n}=n/2(2a+(n-1)d)\] n=number of terms a=first term d=common diff

OpenStudy (anonymous):

so first we make the sequence

OpenStudy (anonymous):

okay :) how do we make the sequence?

OpenStudy (anonymous):

find 1st term 2nd term and so on.by plugging 1,2,3,4...

OpenStudy (anonymous):

plug it into the sum formula??? that you wrote above?

OpenStudy (anonymous):

or into 6n+5?

OpenStudy (anonymous):

no we are jus looking for the terms,so in an=6n+5

OpenStudy (anonymous):

oh okay... so like this? 6(1)+5=11 6(2)+5=17 etc..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so should i list them all out up to 30??

OpenStudy (anonymous):

no the formular makes it easy well see later, so whats a= d= n=

OpenStudy (anonymous):

oh okay! :) haha that would've taken a long time to write out :) so... a=first term which is 11 d=common difference which is 6 n=number of terms which is 30 is that correct?

OpenStudy (anonymous):

yes so look at the sum formular and substitute

OpenStudy (anonymous):

S30=30/2(2(11)+(30-1)(6)) is that right so far?

OpenStudy (anonymous):

and then i get 15(22+29(6)) which = 15(22+174) which = 15(196)=2,940

OpenStudy (anonymous):

so is the answer 2,940 ???? 2,940 is the sum of the first 30 terms of the sequence 6n+5???

OpenStudy (anonymous):

yes so we dont have to write them until 30,exam questions can ask the sum of 300 terms

OpenStudy (anonymous):

oh cool! so i just need to memorize these equations?

OpenStudy (anonymous):

and so 2,940 is my final answer??

OpenStudy (anonymous):

@Jonask the sum formula is the only formula i need to memorize to answer questions such as these yes??

OpenStudy (anonymous):

but sometimnes there is a diff sum form

OpenStudy (anonymous):

oh okay cool!! :) So 2940 is the final answer??? just making sure!! :)

OpenStudy (anonymous):

\[S _{n}=n/2(a+l)\]

OpenStudy (anonymous):

oh okay cool beans! :) Thanks @jonask :) u rock!! :)